### v1

# Sliding Puzzles Gym: A Scalable Benchmark for State Representation in Visual Reinforcement Learning

Abstract
Learning effective visual representations is crucial in open-world environments where agents encounter diverse and unstructured observations. This ability enables agents to extract meaningful information from raw sensory inputs, like pixels, which is essential for generalization across different tasks. However, evaluating representation learning separately from policy learning remains a challenge in most reinforcement learning (RL) benchmarks. To address this, we introduce the Sliding Puzzles Gym (SPGym), a benchmark that extends the classic 15-tile puzzle with variable grid sizes and observation spaces, including large real-world image datasets. SPGym allows scaling the representation learning challenge while keeping the latent environment dynamics and algorithmic problem fixed, providing a targeted assessment of agents’ ability to form compositional and generalizable state representations. Experiments with both model-free and model-based RL algorithms, with and without explicit representation learning components, show that as the representation challenge scales, SPGym effectively distinguishes agents based on their capabilities. Moreover, SPGym reaches difficulty levels where no tested algorithm consistently excels, highlighting key challenges and opportunities for advancing representation learning for decision-making research.
1 Introduction
Learning meaningful representations from raw sensory inputs, such as visual data, is fundamental to reinforcement learning (RL) agents’ ability to generalize across different tasks in complex, open-world environments. In visual RL, agents must process high-dimensional pixel data, extract useful features, and utilize these features for decision-making. This becomes especially crucial as real-world applications demand adaptability to unstructured and diverse observations. However, disentangling representation learning from other learning tasks, such as policy optimization or dynamics modeling, remains a key challenge in RL benchmarks.
Most RL benchmarks, including widely adopted ones like Atari [2], DeepMind Control Suite [16], and DeepMind Lab [1], entangle representation learning with other learning challenges, making it difficult to isolate the agent’s ability to form robust representations. More specialized benchmarks, such as the Distracting Control Suite [15], emphasize robustness to visual noise but focuses on filtering out irrelevant complexity.
We address this gap by introducing the Sliding Puzzles Gym (SPGym), [^1] Code available at https://github.com/bryanoliveira/sliding-puzzles-gym. a scalable RL benchmark designed to test agents’ visual representation learning capabilities. SPGym extends the 15-tile puzzle with variable grid sizes and observation spaces, such as image-based tiles, allowing for a controlled and scalable increase in representation complexity. Crucially, SPGym keeps the underlying environment dynamics and task structure constant, isolating the representation learning challenge.
We evaluate the performance of both model-free and model-based RL agents on SPGym, comparing their ability to generalize across increasingly complex visual input spaces. Our experiments reveal that, as representation learning complexity increases, SPGym effectively differentiates between agents based on their ability to form compositional and generalizable state representations. Specifically, model-based agents such as DreamerV3 [4] outperform model-free ones like PPO [14] in most scenarios, though even these methods struggle with higher levels of visual complexity.
Contributions
We summarize our main contributions as follows:
  • We introduce the Sliding Puzzles Gym (SPGym), a scalable benchmark for assessing representation learning in visual decision-making. SPGym achieves this by gradually increasing the complexity of visual observations without changing the underlying task structure.
  • We provide an experimental evaluation of model-free and model-based RL agents, demonstrating the benchmark’s effectiveness in distinguishing between agents’ representation learning capabilities.
  • We reveal limitations in current RL methods’ ability to handle complex and diverse visual inputs, showcasing the scalability and challenge SPGym offers as a benchmark.
Our experiments show that SPGym offers a challenging testbed for visual representation learning in RL, making it a valuable tool for future research aimed at improving agents’ visual representation capabilities in open-world environments.
2 Related Work
Reinforcement learning benchmarks are essential tools for evaluating agents’ capabilities across diverse tasks. Popular benchmarks like Atari [2], DeepMind Control Suite [16], and DeepMind Lab [1] focus primarily on assessing policy learning and generalization. Yet, these benchmarks inherently combine representation learning with policy optimization and dynamics modeling, making it difficult to isolate the specific impact of representation learning on an agent’s overall performance. Even methods aimed at improving representation learning for RL agents, such as DARLA [5], CURL [8], RAD [9], DrQ [7], CBM [10], and CycAug [11] are typically evaluated within these entangled settings. While these methods improve agents’ ability to learn from raw visual inputs, they do not provide a clear, isolated assessment of their representation learning capabilities.
ProcGen [3] takes a step toward evaluating generalization through diversity by offering procedurally generated levels, challenging agents to adapt to unseen environments. However, like other benchmarks, it does not fully separate representation learning from policy and dynamics learning. Agents must simultaneously learn representations, policies, and environment dynamics, making it difficult to attribute performance improvements solely to better representations. Alternatively, the Distracting Control Suite [15], an extension of DM Control, introduces visual distractions to evaluate agents’ robustness to irrelevant variations. While it allows for parametrizable control over distraction complexity, its primary focus is on agents’ ability to ignore these distractions rather than to extract meaningful information from visually complex observations.
SPGym’s design sets it apart from existing benchmarks by providing a controlled environment where difficulty arises solely from the complexity of visual observations, rather than changes in task dynamics or objectives. This makes SPGym particularly well-suited for comparing RL algorithms that specifically target improvements in representation learning.
3 The Sliding Puzzles Gym
Figure 1: Overview of the Sliding Puzzles Gym (SPGym). The framework extends the 15-tile puzzle by incorporating image-based tiles, allowing scalable representation complexity while maintaining fixed environment dynamics.
The Sliding Puzzles Gym (SPGym) extends the 15-tile puzzle, where players rearrange a shuffled $4\times 4$ grid of numbered tiles into the correct sequence using four actions: UP, DOWN, LEFT, or RIGHT. SPGym generalizes this concept by supporting variable grid sizes, ranging from $2\times 2$ to larger $H\times W$ configurations, and incorporating diverse observation spaces. In this work, we focus on $3\times 3$ grids to emphasize the challenge of visual representation learning using image-based observations.
SPGym’s internal representation is a 2D array of tile indices representing the current grid state. We provide wrappers that convert these states into agent-friendly observations, such as one-hot encodings or, as used in this paper, image overlays. Figure 1 presents the main components of SPGym.
At each time step, the reward is computed as
$r_{t}=\begin{cases}-\frac{\sum_{i,j}|x_{i,j}-x^{*}_{i,j}|+|y_{i,j}-y^{*}_{i,j}|}{D},&\text{if action is valid}\\ -1,&\text{if action is invalid}\\ +10,&\text{if puzzle is solved}\end{cases}\text{,}$ (1)
with
$D=\sum_{i=0}^{H-1}\sum_{j=0}^{W-1}\left[(i,j)\neq(H-1,W-1)\right]\cdot\left(\max(i,H-1-i)+\max(j,W-1-j)\right)\text{.}$ (2)
Here, $(x_{i,j},y_{i,j})$ represents the current position of the tile at index $(i,j)$, $(x^{*}_{i,j},y^{*}_{i,j})$ is its target position, and $D$ is the maximum possible sum of Manhattan distances, excluding the blank tile. Invalid actions, such as attempting to move a tile outside the grid boundaries, result in a penalty of $-1$. For example, in the puzzle shown in Figure 2, the DOWN and RIGHT actions would be invalid. Successfully solving the puzzle rewards the agent with $+10$, and the episode terminates.
SPGym provides two methods for generating the puzzle’s initial state. The first method, used in this paper, generates a random $H\times W$ array and adjusts the parity of the puzzle by swapping the first two tiles if necessary to ensure solvability [6]. The second method, which can support curriculum learning in future work, starts with a solved puzzle and applies a random series of valid moves to generate an initial state, enabling easier starting puzzle configurations.
Wrappers and Observation Modalities
Figure 2: Different observation modalities in SPGym. Each modality presents a unique challenge for representation learning. The four presented observations represent the same latent puzzle state.
SPGym offers flexible wrappers that transform internal states into different modalities, ranging from simple one-hot encodings to complex visual observations, such as real-world images and text (see Figure 2). This flexibility makes SPGym adaptable for exploring how reinforcement learning agents handle diverse observation spaces, enhancing its utility as a benchmark for representation learning. In this paper, however, we focus on visual representation learning, leaving other modalities for future exploration.
For image overlays, we sample $p$ images from a predefined dataset to form a pool for each run. The benchmark is dataset-agnostic, enabling the use of any image dataset, including images generated procedurally via diffusion. At the beginning of each episode, SPGym selects a random image from the run pool, splits it into $H\times W$ indexed patches, and overlays it onto the puzzle. The agent must then solve the puzzle by reconstructing the shuffled image, testing its ability to form compositional visual representations from raw pixel data.
Complexity Scalability
A key feature of our benchmark is its scalable difficulty. SPGym can increase representation learning complexity by keeping the grid size fixed while expanding the pool of images. This approach holds the underlying puzzle dynamics constant, ensuring that the increased difficulty comes solely from the agent’s need to handle a larger variety of visual observations.
SPGym also scales by adjusting grid sizes. Larger grids increase the search space for solving the puzzle, though the core algorithmic problem remains unchanged. [^2] The same algorithm can be used to solve the puzzle, regardless of grid size [12, 17]. Larger grid sizes also increase the complexity of representation learning by splitting images into more patches. Although scaling the grid size helps to evaluate the efficacy of policy-learning algorithms in more challenging settings, we find that a $3\times 3$ grid is sufficient to differentiate the performance of the algorithms tested in this study.
By maintaining fixed environment dynamics and task mechanics across different levels of difficulty, SPGym ensures that performance variations reflect the agent’s ability to learn and generalize visual representations, rather than adapt to changing dynamics. Consequently, SPGym rigorously assesses an agent’s capability to handle increasingly complex visual inputs while preserving consistent underlying behaviors.
4 Experimental Setup
We evaluate the performance of both model-free and model-based reinforcement learning agents using the image-based variation of SPGym. To simulate open-world conditions, we select images from the validation split of the ImageNet-1k [13] dataset. [^3] Available at https://huggingface.co/datasets/ILSVRC/imagenet-1k This choice allows future work to leverage encoders pretrained on the train split, ensuring reproducibility and comparability with our results.
For model-free experiments, we employ Proximal Policy Optimization [14, PPO] agents. In the model-based category, we use agents based on the DreamerV3 [4] architecture, which includes an autoencoder to enhance representation learning. To evaluate how well SPGym differentiates agents based on their representation learning capabilities, we compare the performance of standard PPO and DreamerV3 agents with variants that alter the representation learning process. Specifically, we include a PPO variant with pretrained encoders and a DreamerV3 variant without gradients from the decoder (by setting the decoder’s loss_scale to 00). For the pretrained PPO agents, we use encoders from agents trained on both the same pool and different pools of images. This allows us to assess, respectively, potential improvements in sample efficiency with better encoders and the generalization capabilities of these pretrained models to new visual inputs. We load encoders pretrained on the same pool sizes and leave the exploration of how encoders trained with larger pool sizes generalize to unseen pools for future work. In all configurations, gradients propagate through the entire model.
We measure the agents’ performance primarily through sample efficiency, defined as the number of environment interactions required to reach a predefined threshold of 80% success rate. To prevent excessively long runs, we impose a cap of 10 million environment steps per run, and we apply early stopping if the agents maintain a success rate of 100% for at least 100 episodes. We also limit the agent’s step count per episode to 1,000. Appendix A contains all hyperparameters relevant for result reproduction.
We conduct all experiments using 5 independent seeds and report the average results with corresponding error bars. For Figure 3, error bars represent the 95% confidence interval calculated using bootstrap resampling with 1,000 iterations. In Figure 4, we employ a rolling window average over 100,000 environment steps per run to account for potentially varying log rates across different runs. The 95% confidence interval for this figure is computed using the Z-score method.
Our hardware setup comprises an AMD Ryzen 7 3700X CPU, an NVIDIA RTX 3090 GPU, 64GB of RAM, and 128GB of swap space. Using this configuration, the most time-consuming DreamerV3 runs require approximately 20 hours, while the longest PPO runs complete in around 30 minutes. [^4] We attribute this runtime disparity primarily to DreamerV3’s extensive utilization of swap space for its replay buffer.
5 Results
Figure 3: Average number of steps to solve the puzzle as a function of the pool size. We consider the puzzle solved when the agents reach $80\%$ success rate and limit the maximum number of environment steps per run to 10 million. Error bars represent the $95\%$ confidence interval of a bootstrap with 1,000 resamples from 5 independent seeds.
Figure 4: Success rate as a function of environment steps. The gradual increase in representation complexity affects the sample efficiency of standard PPO and DreamerV3 agents. Error bands indicate the 95% confidence interval, derived from 5 independent seeds and a rolling window of 100,000 environment steps, using the Z-score method.
The results in Figure 3 demonstrate that as the complexity of the image pool increases, model-based agents, particularly the standard DreamerV3, consistently outperform their model-free counterparts. The DreamerV3 agent without the decoder performs noticeably worse, emphasizing the importance of DreamerV3’s autoencoder structure in representation learning. For image pools with two or more images, DreamerV3 maintains a clear advantage over PPO-based agents, which is expected given the absence of dedicated representation learning mechanisms in PPO and the numerous optimizations present in DreamerV3.
Interestingly, we observe that PPO agents with pretrained encoders perform better than those without, even when the pretraining was done on different image pools. This suggests that these encoders capture transferable, task-agnostic features that aid performance. However, as the representation learning challenge increases (i.e., larger image pools), the efficiency of all agents diminished. This trend reflects the scaling difficulty introduced by SPGym, further highlighting its utility as a benchmark for testing agents’ ability to generalize across diverse visual inputs.
Figure 4 illustrates the success rates as a function of environment steps for both standard DreamerV3 and PPO agents, across various image pool sizes. As the pool size increases, both agents exhibit a decline in sample efficiency. Even the DreamerV3 agent, which generally performs well, fails to solve the puzzle when the pool size scales, underscoring the difficulty of scaling representation learning in complex visual environments. This result demonstrates the potential of SPGym to challenge even state-of-the-art agents, revealing limitations in their ability to generalize from raw visual inputs.
Overall, our experiments show that SPGym effectively discriminates between agents based on their ability to form robust and generalizable state representations. The increasing difficulty, driven by visual diversity, highlights key challenges that remain in representation learning for decision-making tasks. These results open up several avenues for future research, particularly in developing agents that can handle more complex and diverse observation spaces.
6 Conclusion
We introduce the Sliding Puzzles Gym (SPGym), a scalable and flexible benchmark designed to evaluate the representation learning capabilities of reinforcement learning (RL) algorithms. By decoupling the complexity of the representation challenge from the underlying task structure, SPGym provides a controlled setting to assess agents’ ability to form compositional and generalizable state representations. Our experiments highlight the effectiveness of approaches like DreamerV3, which outperformed PPO agents in most settings. However, even these advanced methods struggle as the complexity of the puzzle increases, illustrating the potential difficulty of the task. SPGym’s design allows for unlimited scalability, including the potential for generating new image datasets on the fly using diffusion models, further increasing its utility as a benchmark for future research and development of more capable open-world agents.
Limitations
While SPGym presents a versatile framework, this work has several limitations. First, our experiments did not include agents specifically designed for representation learning, such as DARLA, CURL, RAD, DrQ, CBM, and CycAug. Future research should evaluate how SPGym ranks such agents in its challenging settings. Second, the high stochasticity of the environment and agents’ sensitivity to the sampled image pools suggest that more seeds should be used in each experiment to ensure statistical robustness. However, this was impractical in our study due to the prohibitive computational costs.
Future Work
SPGym’s adaptability opens several promising paths for future research. While this paper focuses on visual inputs, the benchmark supports extension to other data modalities, such as text, enabling investigation of how agents handle a variety of input types. Additionally, future work could explore in- and out-of-distribution generalization by training agents and encoders on specific image pools, external datasets, or through unsupervised learning methods, and testing their performance on unseen or novel data classes. Another valuable direction is the integration of curriculum learning into SPGym, where agents begin with simpler, partially solved puzzles and gradually face more difficult configurations. These extensions would deepen our understanding of how RL agents learn and generalize in complex, real-world environments.
Appendix A Hyperparameters
Table 1 lists hyperparameters used across all experiments, unless noted otherwise. For DreamerV3, we adopted hyperparameters from [4], modifying only the decoder loss scale (set to 0) for the version without decoder.
Table 1: Hyperparameters for PPO and DreamerV3
Algorithm Hyperparameter Value
PPO Env steps 10M
Env instances 1024
Optimizer Adam
Learning Rate (LR) 2.5e-4
LR annealing Yes
Adam Epsilon 1e-5
Num. steps 4
Num. epochs 4
Batch size 64
Num. minibatches 4
Gamma 0.99
GAE lambda 0.95
Advantage normalization Yes
Clip coef. 0.1
Clip value loss Yes
Value function coef. 0.5
DreamerV3 Env steps 10M
Env instances 16
Model size 12M
Replay capacity 5e6
Replay ratio 32
Action repeat 1
Learning rate 4e-5
Batch size 16
Batch length 64
Imagination horizon 15
Discount horizon 333
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction accurately capture the paper’s contributions and scope. We clearly present our main claim of introducing SPGym as a new benchmark for state representation in visual reinforcement learning, highlight its key features, and summarize our experimental findings. The results support our claim that SPGym effectively differentiates RL algorithms based on their representation learning capabilities.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Although our results support the main claims, we recognize the need for broader experimentation with different algorithms and an increased number of independent runs for improved statistical significance.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We describe the implementation details of the proposed benchmark in Section 3 and hyperparameters used in our experiments in Appendix A.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general, releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The code for the benchmark and scripts for reproducing our experiments will be provided in Section 1 and as an attatchment to this submission.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide the training and test details in Section 3 and the hyperparameters used in our experiments in Appendix A.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report the error bars as the 95% confidence interval of a bootstrap with 1,000 resamples from 5 independent seeds in Figure 3. For Figure 4 we report the standard deviation from 5 independent seeds and a rolling window of 10,000 environment steps.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide the compute resources needed to reproduce our experiments in Section 4.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The authors have reviewed the NeurIPS Code of Ethics and confirm that the research conducted in the paper conforms to all of its requirements.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We present a foundational benchmark aimed at improving representation learning for decision making systems, which could benefit areas like robotics and autonomous systems. However, risks include misuse in surveillance or unfair decision-making. Mitigation involves ensuring responsible development and deployment.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper does not release any data or models that have a high risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We credit the authors of ImageNet in Section 4.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide implementation details in Section 3 and the documented code for the benchmark and scripts for reproducing our experiments in Section 1 and as an attatchment to this submission.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.